OpenStudy (anonymous):

Two concert tours were by a Jazz band & a rock band. Together they toured 175 cities. The jazz band toured 95 cities more than the rock band. How many cities did the Jazz band visit How many cities did the Rock band visit

OpenStudy (anonymous):

J= number of cities the jazz band toured R=number of cities the rock band toured We know that: 175=J+R and J=95+R Substitute J: 175= (95+R) + R 175= 95 + 2R Subtract 95 from both sides 80= 2R Divide by 2 R=40 Go back to original formulas and plug in R=40, you get J=135 the rock band toured 40 cities the jazz band toured 135 cities

OpenStudy (anonymous):

follow betty solution it's correct. .